use the Laplace transform to solve the given initial value problem:

Questions 8 and 12

8. Y – 2y + 2y = cos t; y(0) = 1, y(0) = 0 9. y – 2y + 2y = e-t; y(0) = 0, y(0) = 1 »Answer Solution 10. y + 2y + y =
Answer #1

s 3+1 SA 6 aiven initial-value problem (INA) is y_24+2y = cost i yo)=1, glosão Applying Laplace pransform, then Lly-24+2yFrom ) D= 2-2B use this in. -2 A+B+ 2-23 5 - 2 => -2 A-CC ~ 나 CD 2AA4 selving and, 2A-4B=2 2AA B - -SB=-2%B= As As 며 내hn A응,.:: ytt) = ŚLacost 72 sint-4et cost +2 etsint). Answer 414} = Ś la cost+2 sint-4etcost +Retsint). (12 civen initial-value pr5 + s (5-1)(3+1) using in Then Y(s) = Ž ša tž si Applying Inverse Laplace Transfoom. Then y(t) = [l È + Ž 4 ) Syd)= {[c15]+[